OpenStudy (anonymous):

56% of what number is 75.6?

OpenStudy (anonymous):

Let the number be x. 56x/100 = 75.6 Solve it to get the answer.

OpenStudy (anonymous):

ans = lambda: 75.6/0.56

OpenStudy (anonymous):

135

OpenStudy (anonymous):

translate in to math 56% means .56 "of" means "times" "what number" make a variable. i choose n "is" is "=" 75.6 is 75.6 translate as \[.56\times n = 75.6\] then solve via \[x=75.6\div .56\]

OpenStudy (anonymous):

42.33600

OpenStudy (anonymous):

^ no sorry 75.6/56*100 = 135